4. Let A= =(-6 ;) -14 6 9 Find the eigenvalues and a basis for each eigenspace.

linear algebra
Answer #1

Given: 4 A = 6 9 Characteristic equation: -1-1 4 IA - 11/20 6 9-1 (1+1)(1-9) – 24 = 0 -> +2 -81-9-24 = 0 » x2-84 -33=0 - x =